# LDMol: Text-to-Molecule Diffusion Model with Structurally Informative Latent Space

### v2

@@ 3.3 Implementation details
The pre-training of the SMILES encoder and the corresponding decoder was done with 10,000,000 general molecules from PubChem (Kim et al., 2023). To ensure enough batch size for negative samples (He et al., 2020), we build a memory queue that stores $Q$ recent input and use them for the encoder training. The SMILES tokenizer vocabulary consists of 300 tokens, which were obtained from the pre-training data SMILES corpus using the BPE algorithm (Gage, 1994). We only used SMILES data that does not exceed a fixed maximum token length $L$.
To train the text-conditioned latent diffusion model, we gathered three existing datasets of text-molecule pair: PubchemSTM curated by Liu et al. (2023a), ChEBI-20 (Edwards et al., 2021), and PCdes (Zeng et al., 2022). Only a train split for each dataset was used for the training, and pairs that appear in the test set for the experiments are additionally removed. We also used 10,000 molecules from ZINC15 (Sterling & Irwin, 2015) without any text descriptions, which helps the model learn the common distribution of molecules. When these unlabeled data were fed into the training model, we used a pre-defined null text which represents the absence of text condition.
The latent diffusion model was trained with the training loss suggested by Dhariwal & Nichol (2021). To take advantage of classifier-free guidance (Ho & Salimans, 2021), we randomly replaced 3% of the given text condition with the null text during the training. The sampling iteration in the inference stage used DDIM-based (Song et al., 2021a) 100 sampling steps with a classifier-free guidance. More detailed training hyperparameters can be found in Appendix A.2.
@@ 4.1 Text-conditioned molecule generation
In this section, we evaluated the trained LDMol’s ability to generate molecules that agree with the given natural language conditions. First, we generated molecules with LDMol using the text captions in the ChEBI-20 test set and compared them with the ground truth. The metrics we’ve used are as follows: SMILES validity, BLEU score (Papineni et al., 2002) and Levenshtein distance between two SMILES, Tanimoto similarity (Bajusz et al., 2015) between two SMILES with three different fingerprints (MACCS, RDK, Morgan), the exact match ratio, and Frechet ChemNet Distance (FCD) (Preuer et al., 2018).
We note that since our diffusion model sampling starts from randomly sampled noise from a prior distribution, the output from deterministic DDIM sampling still has stochasticity. However, with a highly detailed text prompt that can specify a single correct answer, we discovered that LDMol was powerful enough to outperform many autoregressive models with randomly sampled initial points. Also, the performance difference was marginal between starting from random noise and starting from a fixed point (e.g., zero vector). We tested different scales for the classifier-free guidance scale $\omega$ in the sampling process and found $\omega=2.5$ works best (See Section B.1).
Table 1: Benchmark results of text-to-molecule generation on ChEBI-20 test set. The best performance for each metric was written in bold. The “Family” column denotes whether the model is AR(autoregressive model) or DM(diffusion model).
@@
Table 1 contains the performance of LDMol and other baselines for text-to-molecule generation on the ChEBI-20 test set. Including both transformer-based autoregressive models and diffusion-based models, LDMol outperformed the existing models in almost every metric. While few models showed higher validity than ours, they showed a lower agreement between the output and the ground truth, which we insist is a more important role of generative text-to-molecule models. One thing to emphasize is that MolT5large uses the same text encoder with LDMol, yet there’s a significant performance difference between the two models. We believe this is because our constructed latent space is much easier to align with the same textual information, compared to the token sequence used by transformer-based models.
@@
To demonstrate the LDMol’s molecule generalization ability with more broad and general text inputs, we analyzed the generated output with several hand-written prompts. These input prompts were not contained in the training data and were relatively vague and high-level so that many different molecules could satisfy the condition. Figure 4 shows the samples of generated molecules from LDMol with several input prompt examples. We found that LMDol can generate molecules with high validity that follow the various levels of input conditions for specific atoms(a), functional groups(b), molecular substructure(c), compound class(d), and substance names(e). The validity was calculated by the number of valid SMILES over 1,000 generated samples, above 0.9 for most scenarios we tested. Considering that these short, broad, and hand-written text conditions are distinct from the text conditions in the training dataset, we’ve concluded that our model is able to learn the general relation between natural language and molecules.
@@ 4.2 Applications toward downstream tasks
Well-trained diffusion models learned the score function of a data distribution, which enables high applicability to various downstream tasks. The state-of-the-art image diffusion models have shown their versatility in image editing (Meng et al., 2022; Hertz et al., 2023), classification (Li et al., 2023), retrieval (Jin et al., 2023), inverse problems like inpainting and deblurring (Chung et al., 2023), image personalization (Ruiz et al., 2023), etc. To demonstrate LMDol’s potential versatility as a diffusion model, we applied the pre-trained LDMol to the molecule-to-text retrieval and text-guided molecule editing task.
Molecule-to-text retrieval. Our approach to molecule-to-text retrieval is similar to the idea of using a pre-trained diffusion model as a classifier (Li et al., 2023): LDMol takes each candidate text with a query molecule’s noised latent, and retrieves the text that minimizes the noise estimation error $||\hat{\epsilon}_{\theta}-\epsilon||^{2}_{2}$ between the injected noise $\epsilon$ and the predicted noise $\hat{\epsilon}_{\theta}$(see Appendix A.3 for more details). Since this process has randomness due to the stochasticity of $t$ and $\epsilon$, we repeated the same process $n$ times with resampled $t$ and $\epsilon$ and used a mean error to minimize the performance variance.
@@
sentence-level paragraph-level sentence-level paragraph-level
@@
We measured a 64-way in-batch retrieval accuracy of LDMol using two different test sets: PCdes test split and MoMu retrieval dataset curated by Su et al. (2022), where the result with other baseline models are listed in Table 2. Only one randomly selected sentence in each candidate description was used for the retrieval in the “sentence-level” column, and all descriptions were used for the “paragraph-level” column. LDMol achieved a higher performance in all four scenarios compared to the previously presented models and maintained its performance on a relatively out-of-distribution MoMu test set with minimal accuracy drop. LDMol became more accurate as the number of function evaluations increased, and the improvement was more significant at the sentence-level retrieval and out-of-distribution dataset. The actual examples from the retrieval result can be found in Appendix B.2.
Figure 5: (a) Hit ratio of molecule editing by LDMol and MoleculeSTM (Liu et al., 2023a) in eight scenarios. Following MoleculeSTM, each scenario was applied to 200 randomly sampled molecules from ZINC15, and the mean and standard variation on three separate runs were plotted. In the figure, we omitted “This molecule” in front of the actual prompts, and abbreviated “hydrogen bonding” to “H-bond”. (b) Examples of text-guided molecule editing with LDMol. The difference between the source text and the target text, and the corresponding region is colored in purple.
Text-guided molecule editing. We applied a method of Delta Denoising Score (DDS) (Hertz et al., 2023), which was originally suggested for text-guided image editing, to see whether LDMol can be used to optimize a source molecule to match a given text. Using two text prompts that describe the source data $z_{src}$ and the desired target, DDS presents how a text-conditioned diffusion model can modify $z_{src}$ into a new data $z_{tgt}$ that follows the target text prompt (more detailed procedures can be found in Appendix A.3).
We imported a method of DDS on LDMol’s molecule latent to edit a given molecule to match the target text, with several prepared editing prompts that require the model to change certain atoms, substructures, and intrinsic properties from the source molecule. Figure 5-(a) shows that LDMol had comparable performance with a previously suggested text-guided molecule editing model (Liu et al., 2023a), with a higher hit ratio in five out of eight scenarios. Actual case studies with editing sources and results are listed in Figure 5-(b), where the editing results successfully modified the input molecule towards the target prompt with minimal corruption of the unrelated region.
4.3 Ablation studies
We’ve conducted ablation studies for each design choice of the proposed LDMol to analyze and emphasize their role. Each model is pre-trained with the same number of DiT training iterations. The results of the ablation studies are listed in Table 3.
Table 3: Quantitative results of the ablation studies.
models Autoencoder Text-to-molecule generation
@@ 4.3 Ablation studies
LDMol w/o contrastive learning(i.e. naive AE) 1.000 0.019 0.000 58.60
LDMol w/o compression layer 0.964 0.022 0.000 67.93
LDMol w/o stereoisomer hard-negative 0.891 0.939 0.278 0.24
naive AE + KL regularization($\beta=0.001$) 0.999 0.847 0.492 0.34
@@
When we remove the contrastive encoder pre-training objective and construct the molecule latent space with a naive autoencoder, the diffusion model completely fails to learn the latent distribution to generate valid SMILES. A similar problem occurred when we didn’t introduce a compression layer since the dimension of the latent space was too big for the diffusion model to learn. When stereoisomers were not utilized as hard negative samples in the contrastive encoder training, the constructed latent space was not detailed enough to specify the input, which degraded the reconstruction accuracy of the autoencoder.
An interesting case is the comparison against the autoencoder with KL-divergence loss: it has reconstructible latent space and showed a text-to-molecule generation match ratio of 0.492, which outperforms several baseline models in Table 1. This performance shows the ability of diffusion models on the continuous data domain which can be comparable to the autoregressive models. However, its overall metric is still lower than the proposed LDMol, with especially low validity and FCD. We insist that this gap comes from the structurally informative latent space of LDMol which is easier for the diffusion model to learn the correlation between the latent space and the condition.
@@ 5 Conclusion
In this work, we presented a text-conditioned molecule diffusion model LDMol that runs on a chemical latent space reflecting structural information. By introducing the deeply studied paradigm of the latent diffusion model into the molecule domain with minimal modifications, LDMol retains many advanced attributes of diffusion models that enable various applications.
Despite the noticeable performances of LDMol, it still has limitations that can be improved: Since the text captions of our obtained paired data mostly describe structural features, LDMol still struggles to follow some text conditions such as complex biological properties. Nonetheless, we expect that the LDMol’s performance could be improved further with the emergence of richer text-molecule pair data and more powerful text encoders. Moreover, combining physiochemical and biological annotations on top of the structurally informative latent space is a promising future work that can ease the connection between molecules and text conditions.
We believe that our approach could inspire tackling various chemical generation tasks using latent space, not only text-conditioned but also many more desired properties such as biochemical activity. Especially, we expect LDMol to be a starting point to bridge achievements in the state-of-the-art diffusion model into the chemical domain.
@@ A.2 Model hyperparameters and training setup
Detailed hyperparameters on the model architecture are listed in Table 4, with settings on the model training procedure.
@@
A.3 More detailed procedure for LDMol’s application on downstream tasks

### v4

@@ 3.3 Implementation details
The pre-training of the SMILES encoder and the corresponding decoder was done with 10,000,000 general molecules from PubChem (Kim et al., 2023). The SMILES tokenizer vocabulary consists of 300 tokens, which were obtained from the pre-training data SMILES corpus using the BPE algorithm (Gage, 1994). We only used SMILES data that does not exceed a fixed maximum token length $L$. To ensure enough batch size for negative samples (He et al., 2020), we build a memory queue that stores $Q$ recent input and use them for the encoder training. We found that if the training data have a stereoisomer, considering it as “hard-negative” samples and including it in the loss calculation batch helps the encoder training to differentiate different stereoisomers.
To train the text-conditioned latent diffusion model, we gathered three existing datasets of text-molecule pairs: PubchemSTM curated by Liu et al. (2023a), ChEBI-20 (Edwards et al., 2021), and PCdes (Zeng et al., 2022). Only a train split for each dataset was used for the training, and pairs that appeared in the test set for the experiments are additionally removed. We also used 10,000 molecules from ZINC15 (Sterling & Irwin, 2015) without any text descriptions, which helps the model learn the common distribution of molecules. When these unlabeled data were fed into the training model, we used a pre-defined null text for the absence of text condition. We utilized 320,000 training data in total, much smaller than recent transformer-based baselines (Edwards et al., 2022; Pei et al., 2024) with millions of unimodal and multimodal data from various databases.
The latent diffusion model was trained with the training loss suggested by Dhariwal & Nichol (2021). To take advantage of classifier-free guidance (Ho & Salimans, 2021), we randomly replaced 3% of the given text condition with the null text during the training. The sampling iteration in the inference stage used DDIM-based (Song et al., 2021a) 100 sampling steps with a classifier-free guidance. More detailed training hyperparameters can be found in Appendix A.2. The code for LDMol training and text-to-molecule sampling is available at https://github.com/jinhojsk515/LDMol.
@@ 4.1 Text-conditioned molecule generation
Table 1: Benchmark results of text-to-molecule generation on ChEBI-20 and PCDes test set. The best performance for each metric was written in bold. The “Family” column denotes whether the model is AR(autoregressive model) or DM(diffusion model).
@@
PCDes MolT5large AR 0.944 0.692 18.481 0.810 0.741 0.699 0.440 0.70
bioT5 AR 1.000 0.754 15.658 0.797 0.726 0.677 0.455 0.69
bioT5+ AR 0.999 0.677 20.464 0.743 0.615 0.541 0.266 1.09
LDMol DM 0.944 0.857 8.726 0.885 0.817 0.780 0.464 0.32
In this section, we evaluated the trained LDMol’s ability to generate molecules that agree with the given natural language conditions. First, we generated molecules with LDMol using the text captions in the ChEBI-20 test set and compared them with the ground truth. The metrics we’ve used are as follows: SMILES validity, BLEU score (Papineni et al., 2002) and Levenshtein distance between two SMILES, Tanimoto similarity (Bajusz et al., 2015) between two SMILES with three different fingerprints (MACCS, RDK, Morgan), the exact match ratio, and Frechet ChemNet Distance (FCD) (Preuer et al., 2018). We tested different scales for the classifier-free guidance scale $\omega$ in the sampling process and found $\omega=2.5$ works best (See Section B.2).
Table 1 contains the performance of LDMol and other baselines for text-to-molecule generation on the ChEBI-20 and PCDes test set. Including both autoregressive models and diffusion-based models, LDMol outperformed the existing models in almost every metric. While few models showed higher validity than ours, they showed a lower agreement between the output and the ground truth, which we insist is a more important role of generative text-to-molecule models. Also, MolT5large uses the same text encoder as LDMol, yet there’s a significant performance difference between the two models. We believe this is because our continuous and structure-aware latent space is much easier to learn and align with the same textual information, compared to the raw token sequence for transformer-based models.
@@
To demonstrate the LDMol’s molecule generalization ability with more broad and general text inputs, we analyzed the generated output with several hand-written prompts. These input prompts were not contained in the training data and were relatively vague and high-level so that many different molecules could satisfy the condition. Figure 4 shows the samples of generated molecules from LDMol with several input prompt examples. We found that LMDol can generate molecules with high validity that follow the various levels of input conditions for specific atoms(a), compound class(b), molecular substructure(c), functional groups(d), and substance names(e). The validity was calculated by the number of valid SMILES over 1,000 generated samples, above 0.9 for most scenarios we tested. Considering that these short, broad, and hand-written text conditions are distinct from the text conditions in the training dataset, we’ve concluded that our model is able to learn the general relation between natural language and molecules. We conducted quantitative analyses on the case studies and additional examples with hand-written prompts, which can be found in Appendix B.3.
@@ 4.2 Applications toward downstream tasks
Well-trained diffusion models learned the score function of a data distribution, which enables high applicability to various downstream tasks. The state-of-the-art image diffusion models have shown their versatility in image editing (Meng et al., 2022; Hertz et al., 2023), classification (Li et al., 2023), retrieval (Jin et al., 2023), inverse problems like inpainting and deblurring (Chung et al., 2023), image personalization (Ruiz et al., 2023), etc. To demonstrate LMDol’s potential versatility as a diffusion model, we applied the pre-trained LDMol to the molecule-to-text retrieval and text-guided molecule editing. See Appendix A.3 for a more detailed procedure for each downstream task.
Molecule-to-text retrieval. Our approach to molecule-to-text retrieval is similar to the idea of using a pre-trained diffusion model as a classifier (Li et al., 2023): LDMol takes each candidate text with a query molecule’s noised latent, and retrieves the text that minimizes the noise estimation error $||\hat{\epsilon}_{\theta}-\epsilon||^{2}_{2}$ between the injected noise $\epsilon$ and the predicted noise $\hat{\epsilon}_{\theta}$. Since this process has randomness due to the stochasticity of $t$ and $\epsilon$, we repeated the same process $n$ times with resampled $t$ and $\epsilon$ and used a mean error to minimize the performance variance.
@@
sentence paragraph sentence paragraph
@@
We measured a 64-way in-batch retrieval accuracy of LDMol using two different test sets: PCdes test split and MoMu retrieval dataset curated by Su et al. (2022), where the result with other baseline models are listed in Table 2. Only one randomly selected sentence in each candidate description was used for the retrieval in the “sentence” column, and all descriptions were used for the “paragraph” column. LDMol achieved a higher performance in all four scenarios compared to the previously presented models and maintained its performance on a relatively out-of-distribution MoMu test set with minimal accuracy drop. LDMol became more accurate as the number of function evaluations increased, and the improvement was more significant at the sentence-level retrieval and out-of-distribution dataset. The actual examples from the retrieval result can be found in Appendix B.4.
Figure 5: Hit ratio of molecule editing by LDMol and MoleculeSTM (Liu et al., 2023a) in eight scenarios. In the figure, we omitted “This molecule” in front of the actual prompts, and abbreviated “hydrogen bonding” to “H-bond”.
Text-guided molecule editing. We applied a method of Delta Denoising Score (DDS) (Hertz et al., 2023), which was originally suggested for text-guided image editing, to see whether LDMol can be used to optimize a source molecule to match a given text. Using two text prompts that describe the source data $z_{src}$ and the desired target, DDS presents how a text-conditioned diffusion model can modify $z_{src}$ into a new data $z_{tgt}$ that follows the target text prompt.
We imported a method of DDS on LDMol’s molecule latent to edit a given molecule to match the target text, with several prepared editing prompts that require the model to change certain atoms, substructures, and intrinsic properties from the source molecule. Figure 5 shows that LDMol had comparable performance with a previously suggested text-guided molecule editing model (Liu et al., 2023a), with a higher hit ratio in five out of eight scenarios. Several editing examples with hand-written scenarios are shown in Appendix B.5.
4.3 Effectiveness of the suggested latent space
We’ve conducted an ablation study in Table 3 that compares LDMol with latent diffusion models trained on naively constructed latent space, to emphasize the benefit of the suggested encoder training. Each model is pre-trained with the same number of DiT training iterations. We’ve also performed an ablation study on more detailed design choices of LDMol, which can be found in Appendix B.6.
Table 3: Quantitative results of the ablation study. The best performance for each metric is written in bold.
models Autoencoder ChEBI20 generation
@@ 4.3 Effectiveness of the suggested latent space
LDMol w/o contrastive learning 1.000 0.019 0.000 58.60
LDMol w/ $\beta$-VAE ($\beta$=0.001) 0.999 0.847 0.492 0.34
@@
When we remove the contrastive encoder pre-training objective and construct the molecule latent space with a naive autoencoder, the diffusion model completely fails to learn the latent distribution to generate valid SMILES. On the other hand, a $\beta$-VAE with KL-divergence regularization has reconstructible latent space and showed a text-to-molecule generation match ratio of 0.492, which already outperforms the previous diffusion model TGM-DLM and several autoregressive models in Table 1. This demonstrates the necessity of diffusion models in the continuous data domain, with their potential to be successful on discrete molecule data comparable to the autoregressive models. Nonetheless, its overall metric is still worse than the proposed LDMol, with notably low validity and FCD. We insist that this gap comes from the structurally informative latent space of LDMol which is easier for the diffusion model to learn the correlation between the latent space and the condition.
@@ 5 Conclusion
In this work, we presented a text-to-molecule diffusion model LDMol that runs on a chemical latent space reflecting structural information. By introducing the deeply studied paradigm of the latent diffusion model with carefully designed latent encoder, LDMol retains many advanced attributes of diffusion models that enable various applications.
Despite the noticeable performances of LDMol, it still has limitations that can be improved, as LDMol still often struggles to follow some text conditions such as complex biological properties. Nonetheless, we expect that the LDMol’s performance could be improved further with the emergence of richer text-molecule pair data and more powerful text encoders. Moreover, combining physiochemical and biological annotations on top of the structurally informative latent space is a promising future work that can ease the connection between molecules and text conditions.
We believe that our approach could inspire tackling various chemical generation tasks using latent space, not only text-conditioned but also many more desired properties, such as biochemical activity. Especially, we expect LDMol to be a starting point to bridge achievements in the state-of-the-art diffusion model into the chemical domain.
Acknowledgments
This work was supported by the National Research Foundation of Korea under Grant No. RS-2024-00336454, and the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2025-02304967, AI Star Fellowship(KAIST)).
Impact Statement
This paper presents work whose goal is to advance the field of machine learning and its chemical applications. There are many potential societal consequences of our work, including the possibility of the molecular design for harmful or inappropriate properties.
@@ A.2 Model hyperparameters and training setup
The LDMol encoder and decoder consist of 12 transformer layers of BERTbase, where the decoder has a causal mask in its self-attention layers and includes a cross-attention layer after each self-attention layer to receive latent information. Detailed hyperparameters on the model architecture are listed in Table 4, with settings on the model training procedure. Here, we provide the results of several ablation studies to support our selection of the hyperparameters.
@@
training resources 8 NVIDIA A100(VRAM:40GB)
Figure 7: The encoder loss convergence with different temperature parameter $\tau$.
Methods Recon. Acc.
Contrastive learning on compressed latent with $d_{z}$=64 0.084
Contrastive learning on $d_{enc}$=1024 compression with $d_{z}$=32 0.948
compression with $d_{z}$=64 0.980
compression with $d_{z}$=128 0.989
Table 5: SMILES reconstruction accuracy of various trained autoencoders, with 1,000 unseen SMILES.
It is known that lower temperature $\tau$ in contrastive learning penalizes hard negatives more strongly, making the learned feature more sensitive to fine-grain details (Wang & Liu, 2021). We considered this as a desirable property for our latent space and used a small tau of 0.07. When we used too big $\tau$ of 0.15 as shown in Figure 7, it reduced the model’s ability to distinguish different molecules and made the training loss converge to a much higher value.
Table 5 lists the LDMol autoencoder’s SMILES reconstruction accuracy in various autoencoder training strategies. When we apply contrastive loss directly into the compressed latent domain, the encoder fails to capture informative features, makes the decoder couldn’t reconstruct the input molecule. In the scenario of adding linear compression after contrastive training with $d_{enc}$=1024, we observed an error rate of more than 5% for the compression size of $d_{z}$=32. Compression with $d_{z}$=128 slightly increased the reconstruction accuracy compared to $d_{z}=64$, but the training time for the subsequent diffusion model rapidly increased. Considering that the failed 2% for the current model were mostly very long molecules, we concluded that $d_{z}=64$ is sufficient for our model.
A.3 LDMol’s application on downstream tasks
Algorithm 1 Molecule-to-Text Retrieval with LDMol
0:  $z,\mathcal{C}=\{c_{i}\}^{B}_{i=1},n\in\mathbb{N}^{+}$
1:  Initialize $\text{Initialize }\texttt{Errors}[c_{i}]=0\textbf{ for }i=1\textbf{ to }B$
2:  for iter=1 to $\texttt{iter}=1\textbf{ to }n$ do
3:     $t\sim U[0,T],\epsilon\sim\mathcal{N}(0,I)$
4:     $z_{t}=\sqrt{\overline{\alpha}_{t}}z+\sqrt{1-\overline{\alpha}_{t}}\epsilon$
5:     for $i=1\textbf{ to }B$ do
6:        $\texttt{Errors}[c_{i}]\mathrel{+}=||\hat{\epsilon}_{\theta}(z_{t},t,c_{i})-\epsilon||_{2}^{2}$
7:     end for
8:  end for
9:  Return $\text{argmin}_{c_{i}\in\mathcal{C}}\texttt{Errors}[c_{i}]$
Algorithm 2 Text-guided Molecule Editing with LDMol
0:  $z_{src},c_{src},c_{tgt},N\in\mathbb{N}^{+},\gamma>0,\omega\geq 1,\mathcal{D}$
1:  Initialize $\text{Initialize }z_{tgt}=z_{src}$
2:  for iter=1 to $\texttt{iter}=1\textbf{ to }N$ do
3:     $t\sim U[0,T],\epsilon\sim\mathcal{N}(0,I)$
4:     $z_{t,src}=\sqrt{\overline{\alpha}_{t}}z_{src}+\sqrt{1-\overline{\alpha}_{t}}\epsilon$
5:     $z_{t,tgt}=\sqrt{\overline{\alpha}_{t}}z_{tgt}+\sqrt{1-\overline{\alpha}_{t}}\epsilon$
6:     $\epsilon_{\theta,src}^{\omega}=(1-\omega)\epsilon_{\theta}(z_{t,src},t,\varnothing)+\omega\epsilon_{\theta}(z_{t,src},t,c_{src})$
7:     $\epsilon_{\theta,tgt}^{\omega}=(1-\omega)\epsilon_{\theta}(z_{t,tgt},t,\varnothing)+\omega\epsilon_{\theta}(z_{t,tgt},t,c_{tgt})$
8:     $z_{tgt}=z_{tgt}-\gamma(\epsilon_{\theta,tgt}^{\omega}-\epsilon_{\theta,src}^{\omega})$
9:  end for
10:  Return $\mathcal{D}(z_{tgt})$